# Exact Model Reduction for Continuous-Time Open Quantum Dynamics

### v2

@@ VIII Acknowledgments
Appendix A Technical proofs and supplementary results
@@
Let then consider the right-hand side of the statement, i.e. $e^{\mathcal{L}t}(\rho_{0})$. Applying the first two facts we just recalled and the definition of the exponential super operator we obtain:
@@
$e^{\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}}t}\Pi_{\mathscr{V}}(\rho_{0})$ $=e^{\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}}t}\underbrace{\Pi_{\mathscr{V}}\Pi_{\mathscr{R}}}_{\Pi_{\mathscr{R}}}(\rho_{0})=\sum_{k=0}^{+\infty}\frac{t^{k}}{k!}(\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}})^{k}\Pi_{\mathscr{R}}(\rho_{0})=\sum_{k=0}^{+\infty}\frac{t^{k}}{k!}(\Pi_{\mathscr{R}}\mathcal{L}\Pi_{\mathscr{R}})^{k}\Pi_{\mathscr{R}}(\rho_{0})=e^{\Pi_{\mathscr{R}}\mathcal{L}\Pi_{\mathscr{R}}t}\Pi_{\mathscr{R}}(\rho_{0}),$
@@
We will prove the proposition by proving the equivalent relation $\Phi^{\prime}=({\rm alg}(\Phi))^{\prime}$, which follows from the Von Neumann bi-commutant theorem, see e.g., [74, Sec. I.9.1]. Then, from the definition of ${\rm alg}(\Phi)$, it follows that $\Phi\subseteq{\rm alg}(\Phi)$, which in turn implies that $\Phi^{\prime}\supseteq({\rm alg}(\Phi))^{\prime}$.
@@ B.1 Single-axis coupling Hamiltonian
Before discussing the more general central-spin model of Sec. VI.1, it is worth considering the simpler yet relevant case where the system-bath interaction Hamiltonian involves a single coupling operator, which also enters the free bath Hamiltonian, say,
@@
since $\sum_{\ell}\left|\varphi_{\ell}\middle>\middle<\varphi_{\ell}\right|=\mathbb{1}_{B}$. In order to practically compute the observable-based MR described in Sec. IV.2, we need to first compute the generators of $\mathscr{N}^{\perp}$, ${\rm ad}^{n}_{H_{SB}}(\sigma_{q}^{(1)})$, with $n=0,1,\dots,2^{2N}-1$. It is then easy to verify that ${\rm ad}^{n}_{H_{SB}}(\sigma_{q}^{(1)})=\sum_{\ell}{\rm ad}^{n}_{H_{S,\ell}}(\sigma_{q})\otimes\left|\varphi_{\ell}\middle>\middle<\varphi_{\ell}\right|$. This fact is already sufficient to prove that the coherences of the bath are not necessary to reproduce the evolution of $\rho_{S}$ for any initial condition $\rho_{0}\in\mathfrak{D}(\mathcal{H})$ and thus the bath can be reduced to a classical Markov model of size $2^{N_{B}}$. However, further reduction is possible in this case.
@@ B.2 Generic XYZ coupling Hamilltonian
We now derive the reduced model presented in Sec. VI.1.1, that is, we consider system-bath model Hamiltonians of the form:
@@
where $K_{n,m}$ are tensor swap permutation matrices, that is, such that $K_{n,m}(A\otimes B)K_{n,m}^{\dagger}=B\otimes A$ for all $A\in\mathbb{C}^{n\times n}$ and $B\in\mathbb{C}^{m\times m}$. Note that the tensor swap matrices are necessary for technical reasons: specifically are included due to the fact that the tensor product is distributive with respect to the diagonal sum $\oplus$ only on the left term of the product, i.e., $\oplus_{k}(A\otimes B_{k})\neq A\otimes(\oplus_{k}B_{k})$, while $\oplus_{k}(A_{k}\otimes B)=(\oplus_{k}A_{k})\otimes B$. The resulting unitary $U$ is such that
@@ Appendix C Weak symmetry in boundary-driven XXZ models
The XXZ boundary-driven model discussed in Sec. VI.2 has been studied in multiple works, including [61, 128]. Let us consider the continuous unitary group $\mathscr{G}_{\varphi}=e^{-i\varphi M}$ generated by the total magnetization operator. Ref. [61] states that $\mathscr{G}_{\varphi}$ is a strong symmetry group for a similar model to the one considered here (same Hamiltonian, different noise operators). However, it was not pointed out that it is also a weak symmetry for the same model we consider. Moreover, [128] states that $\mathscr{G}_{\varphi}$ is a symmetry group for the non-equilibrium SS (see Footnote 5 therein); however, this does not imply that it is a symmetry for the dynamics. Here, we explicitly prove that $\mathscr{G}_{\varphi}$ is a group of weak symmetries for the boundary-driven XXZ chain we consider. We note that similar calculations have been carried out for a XXZ Hamiltonian in [42] (see in particular Sec. II.B.2).
First, since the operators in the set $\{\sigma_{z}^{(j)}\}_{j=1}^{N}$ are mutually commuting, we may write $U_{\varphi}=\prod_{j=1}^{N}U_{\varphi,j}$ where, by using Pauli algebra, we have $U_{\varphi,j}=e^{-i\varphi\sigma_{z}^{(j)}}=\cos(\varphi)\mathbb{1}_{2^{N}}-i\sin(\varphi)\sigma_{z}^{(j)}$.
@@
Let us consider $\sigma_{+}^{(k)}$ and $\sigma_{-}^{(k)}$, with $k=1,\dots,N$. Then $U_{\varphi}\sigma_{+}^{(k)}=e^{-2i\varphi}\sigma_{+}^{(k)}U_{\varphi}$ and $U_{\varphi}\sigma_{-}^{(k)}=e^{2i\varphi}\sigma_{-}^{(k)}U_{\varphi}$.
@@
$U_{\varphi}\sigma_{+}^{(k)}$ $=\prod_{j=1}^{N}U_{\varphi,j}\sigma_{+}^{(k)}=U_{\varphi,k}\sigma_{+}^{(k)}\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}=(\cos(\varphi)\mathbb{1}_{2^{N}}-i\sin(\varphi)\sigma_{z}^{(k)})(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})\prod_{j=2}^{N}U_{\varphi,j}$
$=[\cos(\varphi)(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})-i\sin(\varphi)(\sigma_{z}^{(k)}\sigma_{x}^{(k)}+i\sigma_{z}^{(k)}\sigma_{y}^{(k)})]\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}$
$=[\cos(\varphi)(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})-i\sin(\varphi)({i\sigma_{y}^{(k)}}+{\sigma_{x}^{(k)}})]\prod_{j=2}^{N}U_{\varphi,j}=e^{-i\varphi}\sigma_{+}^{(k)}\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}$
$=e^{-i\varphi}\sigma_{+}^{(k)}U_{\varphi,1}^{\dagger}\underbrace{U_{\varphi,1}\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}}_{U_{\varphi}}=e^{-i\varphi}(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})(\cos(\varphi)\mathbb{1}_{2^{N}}+i\sin(\varphi)\sigma_{z}^{(k)})U_{\varphi}$
$=e^{-i\varphi}[\cos(\varphi)(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})-i\sin(\varphi)({i\sigma_{y}^{(k)}}+{\sigma_{x}^{(k)}})]U_{\varphi}=e^{-2i\varphi}\sigma_{+}^{(k)}U_{\varphi}.$
One similarly finds $U_{\varphi}\sigma_{-}^{(k)}=e^{2i\varphi}\sigma_{-}^{(k)}U_{\varphi}$. ∎
@@
$U_{\varphi,j}\sigma_{x}^{(j)}U_{\varphi,j}^{\dagger}$ $=(\cos(\varphi)\mathbb{1}_{2^{N}}-i\sin(\varphi)\sigma_{z}^{(j)})\sigma_{x}^{(j)}(\cos(\varphi)\mathbb{1}_{2^{N}}+i\sin(\varphi)\sigma_{z}^{(j)})$
$=(\cos(\varphi)\mathbb{1}_{2^{N}}-i\sin(\varphi)\sigma_{z}^{(j)})(\cos(\varphi)\sigma_{x}^{(j)}+i\sin(\varphi)\sigma_{x}^{(j)}\sigma_{z}^{(j)})$
$=\cos^{2}(\varphi)\sigma_{x}^{(j)}+i\cos(\varphi)\sin(\varphi)\sigma_{x}^{(j)}\sigma_{z}^{(j)}-i\cos(\varphi)\sin(\varphi)\sigma_{z}^{(j)}\sigma_{x}^{(j)}+\sin^{2}(\varphi)\sigma_{z}^{(j)}\sigma_{x}^{(j)}\sigma_{z}^{(j)})$
$=[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{x}^{(j)}+2\cos(\varphi)\sin(\varphi)\sigma_{y}^{(j)}.$
Similarly, one obtains $U_{\varphi,j}\sigma_{y}^{(j)}U_{\varphi,j}^{\dagger}=[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{y}^{(j)}-2\cos(\varphi)\sin(\varphi)\sigma_{x}^{(j)}$. Combining these results for $\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}$ and $\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}$, one obtains
$U_{\varphi}\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}U_{\varphi}^{\dagger}$ =$=\left\{[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{x}^{(j)}+2\cos(\varphi)\sin(\varphi)\sigma_{y}^{(j)}\right\}\left\{[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{x}^{(j+1)}+2\cos(\varphi)\sin(\varphi)\sigma_{y}^{(j+1)}\right\},$
$U_{\varphi}\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}U_{\varphi}^{\dagger}$ =$=\left\{[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{y}^{(j)}-2\cos(\varphi)\sin(\varphi)\sigma_{x}^{(j)}\right\}\left\{[\cos^{2}(\varphi)-\sin^{2}(\varphi)]\sigma_{y}^{(j+1)}-2\cos(\varphi)\sin(\varphi)\sigma_{x}^{(j+1)}\right\},$
which, using the identity $[\cos^{2}(\varphi)-\sin^{2}(\varphi)]^{2}=1-4\cos^{2}(\varphi)\sin^{2}(\varphi)$, leads to their sum $U_{\varphi}\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}U_{\varphi}^{\dagger}+U_{\varphi}\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}U_{\varphi}^{\dagger}=\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}+\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}$. ∎

### v3

@@ 8 Acknowledgments
Appendix A Supplementary theory results
A.1 Technical proofs
@@
Let then consider the right-hand side of the statement, i.e., $e^{\mathcal{L}t}(\rho_{0})$. Applying the first two facts we just recalled and the definition of the exponential super operator we obtain:
@@
$e^{\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}}t}\Pi_{\mathscr{V}}(\rho_{0})$ $=e^{\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}}t}\underbrace{\Pi_{\mathscr{V}}\Pi_{\mathscr{R}}}_{\Pi_{\mathscr{R}}}(\rho_{0})=\sum_{k=0}^{+\infty}\frac{t^{k}}{k!}(\Pi_{\mathscr{V}}\mathcal{L}\Pi_{\mathscr{V}})^{k}\Pi_{\mathscr{R}}(\rho_{0})$
$=\sum_{k=0}^{+\infty}\frac{t^{k}}{k!}(\Pi_{\mathscr{R}}\mathcal{L}\Pi_{\mathscr{R}})^{k}\Pi_{\mathscr{R}}(\rho_{0})=e^{\Pi_{\mathscr{R}}\mathcal{L}\Pi_{\mathscr{R}}t}\Pi_{\mathscr{R}}(\rho_{0}),$
@@
We will prove the proposition by proving the equivalent relation $\Phi^{\prime}=({\rm alg}(\Phi))^{\prime}$, which follows from the Von Neumann bi-commutant theorem, see e.g., [78, Sec. I.9.1]. Then, from the definition of ${\rm alg}(\Phi)$, it follows that $\Phi\subseteq{\rm alg}(\Phi)$, which in turn implies that $\Phi^{\prime}\supseteq({\rm alg}(\Phi))^{\prime}$.
@@
A.2 Algorithm complexity
Here, we discuss the implementation and the complexity of the proposed MR algorithms. We focus first on the observable-based linear and quantum MR algorithms: the linear one is covered by the first three steps described below, while to obtain a quantum reduction one has to further proceed by computing minimal algebras and their block decompositions. As in the main text, let $\dim(\mathcal{H})=n<\infty$ and consider a set of $m$ linearly independent observables of interest.
  1. 1.
    In order to compute the generators of $\mathscr{N}^{\perp}$, using Eq. (26), we need to perform the product between a $n^{2}\times n^{2}$ matrix and a $n^{2}$ vector, a number of times equal to $n^{2}-1$ for each observable. This implies a worst-case complexity of $O(mn^{8})$.
  2. 2.
    In order to find an orthogonal basis for $\mathscr{N}^{\perp},$ we can use Gram-Schmidt orthogonalization procedure on the $mn^{2}$ vectors of dimension $n^{2}$, giving a complexity for this step of $O(m^{2}n^{6})$, [146].
  3. 3.
    In practice, the two above steps can be combined to obtain a more efficient algorithm in the more relevant case $\mathscr{N}^{\perp}\subsetneq\mathfrak{B}(\mathcal{H}).$ This follows from the fact that, if we define
    $\mathscr{N}^{\perp}_{f}\equiv{\rm span}\{\mathcal{L}^{{\dagger},k}(O_{j}),k=0,\dots,f,\,\forall j\},$
    we have $\mathscr{N}^{\perp}_{f-1}\subseteq\mathscr{N}^{\perp}_{f}\subseteq\mathscr{N}^{\perp}_{f+1}\subseteq{\mathscr{N}^{\perp}}$. Furthermore, one can prove that whenever $\dim(\mathscr{N}^{\perp}_{f-1})=\dim(\mathscr{N}^{\perp}_{f})$ for some $f$, then $\dim(\mathscr{N}^{\perp}_{f})=\dim(\mathscr{N}^{\perp}_{q})$ for any $q\geq f$, [45]. This implies that, instead of computing $mn^{2}$ vectors and then performing the orthogonalization procedure on all of them one can, at each step, compute $m$ new vectors (which requires a complexity of $O(mn^{6})$) and perform the orthogonalization only on these new vectors (which requires a complexity of $O(mn^{2})$), discarding all the vectors that are linearly dependent while keeping the others. The algorithm stops when a new iteration does not add any new vector, thus repeating the process $f$ times. This reduces the complexity of these two steps from $O(mn^{8})$ to $O(fmn^{6})$. Clearly, $f$ cannot be known a priori and could be as large as $n^{2}$, but this proves to be a more efficient implementation in common cases, as one usually stops for $f<n^{2}$.
  4. 4.
    For quantum MR, it is also necessary to compute (4a) the observable algebra $\mathscr{O}$ and (4b) its Wedderburn decomposition.
    • (4a)
      Since the Wedderburn decomposition of $\mathscr{O}$ and its commutant $\mathscr{O}^{\prime}$ are complementary, one can decide to compute one or the other. To compute the algebra $\mathscr{O}$ from $\mathscr{N}^{\perp}$, one can adapt the standard algorithms for the closure of a Lie algebra, upon substituting the binary operation [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] with the matrix product. Such an approach has a complexity of approximately $O(n^{8})$ [147], although more efficient numerical algorithms are being actively investigated [148]. On the other hand, to compute $\mathscr{O}^{\prime}$, assuming that $\dim(\mathscr{N}^{\perp})=d$, we have to compute the null space of a matrix of size $dn^{2}\times n^{2}$, which implies a complexity of $O(dn^{6})$ using standard SVD techniques [149].
    • (4b)
      The computation of the Wedderburn decomposition requires the computation of the SVD of generic operators in $\mathscr{O}$ (or $\mathscr{O}^{\prime}$) and thus has a complexity of $O(n^{3})$. We also refer the interested reader to [150, 151, 82] for more in-depth discussion.
A complete algorithm for linear ME consists of first performing the linear observable-based MR, followed by the reachable-based MR – without the need to iterate the process; see [51, 45]. Accordingly, based on the above estimates (first three steps), the linear MR algorithm has an overall complexity of $O(n^{8})$.
For quantum MR, the overall complexity of the Algorithm 1 still combines the complexities of the reachable-based and observable-based MR. Unlike the linear reduction, however, the (finite) number of of times that the two subroutines need to be repeated remains an open problem at this time; see [51]. Based on the above estimates, a single run of the algorithm proposed in Sec. 4.2 for observable-based quantum MR has a complexity of the order of $O(n^{8})$. Since the reachable-based quantum MR algorithm proposed in Sec.  3.3 is virtually identical from a computational complexity point of view, it also has a complexity of the order of $O(n^{8})$. Regarding the required number of iterations, a worst-case upper bound is clearly provided by the dimension of the initial operator space ($n^{2}$). However, at each iteration, the size of the operators we manipulate may be reduced, making this bound very conservative. For instance, it is worth stressing that, in all the examples we considered, we never had to iterate more than 3 times.
For all the examples included in this paper, we were able to test the proposed numerical MR algorithm on a commercial laptop for systems up to $N=6$ qubits (although the code can certainly be optimized further). Despite this being a modest number of qubits, it allowed us validate the minimality of the algebra we obtained with symmetry-based considerations, as shown in Fig. 2 and 8. Leveraging the analytical description of the algebra given by symmetry-based considerations, we could then extend the results and compute the reduced model for larger system size, as shown e.g., in Fig. 3 and 4.
@@ B.1 Single-axis coupling Hamiltonian
Before discussing the more general central-spin model of Sec. 6.1, it is worth considering the simpler yet relevant case where the system-bath interaction Hamiltonian involves a single coupling operator, which also enters the free bath Hamiltonian, say,
@@
since $\sum_{\ell}\left|\varphi_{\ell}\middle>\middle<\varphi_{\ell}\right|=\mathbb{1}_{B}$. In order to practically compute the observable-based MR described in Sec. 4.2, we need to first compute the generators of $\mathscr{N}^{\perp}$, ${\rm ad}^{n}_{H_{SB}}(\sigma_{q}^{(1)})$, with $n=0,1,\dots,2^{2N}-1$. It is then easy to verify that ${\rm ad}^{n}_{H_{SB}}(\sigma_{q}^{(1)})=\sum_{\ell}{\rm ad}^{n}_{H_{S,\ell}}(\sigma_{q})\otimes\left|\varphi_{\ell}\middle>\middle<\varphi_{\ell}\right|$. This fact is already sufficient to prove that the coherences of the bath are not necessary to reproduce the evolution of $\rho_{S}$ for any initial condition $\rho_{0}\in\mathfrak{D}(\mathcal{H})$ and thus the bath can be reduced to a classical Markov model of size $2^{N_{B}}$. However, further reduction is possible in this case.
@@ B.2 Generic XYZ coupling Hamilltonian
We now derive the reduced model presented in Sec. 6.1.1, that is, we consider system-bath model Hamiltonians of the form:
@@
where $K_{n,m}$ are tensor swap permutation matrices, that is, such that $K_{n,m}(A\otimes B)K_{n,m}^{\dagger}=B\otimes A$ for all $A\in\mathbb{C}^{n\times n}$ and $B\in\mathbb{C}^{m\times m}$. Note that the tensor swap matrices are necessary for technical reasons: specifically, they are included due to the fact that the tensor product is distributive with respect to the diagonal sum $\oplus$ only on the left term of the product, i.e., $\oplus_{k}(A\otimes B_{k})\neq A\otimes(\oplus_{k}B_{k})$, while $\oplus_{k}(A_{k}\otimes B)=(\oplus_{k}A_{k})\otimes B$. The resulting unitary $U$ is such that
@@ Appendix C Weak symmetry in boundary-driven XXZ models
The XXZ boundary-driven model discussed in Sec. 6.2 has been studied in multiple works, including [65, 130]. Let us consider the continuous unitary group $\mathscr{G}_{\varphi}=e^{-i\varphi M}$ generated by the total magnetization operator. Ref. [65] states that $\mathscr{G}_{\varphi}$ is a strong symmetry group for a similar model to the one considered here (same Hamiltonian, different noise operators). However, it was not pointed out that it is also a weak symmetry for the same model we consider. Moreover, [130] states that $\mathscr{G}_{\varphi}$ is a symmetry group for the non-equilibrium SS (see Footnote 5 therein); however, this does not imply that it is a symmetry for the dynamics. Here, we explicitly prove that $\mathscr{G}_{\varphi}$ is a group of weak symmetries for the boundary-driven XXZ chain we consider. We note that similar calculations have been carried out for a XXZ Hamiltonian in [42] (see in particular Sec. II.B.2).
First, since the operators in the set $\{\sigma_{z}^{(j)}\}_{j=1}^{N}$ are mutually commuting, we may write $U_{\varphi}=\prod_{j=1}^{N}U_{\varphi,j}$ where, by using Pauli algebra, we have $U_{\varphi,j}=e^{-i\varphi/2\sigma_{z}^{(j)}}=\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}-i\sin(\frac{\varphi}{2})\sigma_{z}^{(j)}$.
@@
Consider $\sigma_{+}^{(k)}$ and $\sigma_{-}^{(k)}$, with $k=1,\dots,N$. Then, $U_{\varphi}\sigma_{+}^{(k)}=e^{-i\varphi}\sigma_{+}^{(k)}U_{\varphi}$ and $U_{\varphi}\sigma_{-}^{(k)}=e^{i\varphi}\sigma_{-}^{(k)}U_{\varphi}$.
@@
$U_{\varphi}\sigma_{+}^{(k)}$ $=\prod_{j=1}^{N}U_{\varphi,j}\sigma_{+}^{(k)}=U_{\varphi,k}\sigma_{+}^{(k)}\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}=(\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}-i\sin(\frac{\varphi}{2})\sigma_{z}^{(k)})\sigma_{+}^{(k)}\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}$
$=[\cos(\frac{\varphi}{2})\sigma_{+}^{(k)}-i\sin(\frac{\varphi}{2})(\underbrace{\sigma_{z}^{(k)}\sigma_{x}^{(k)}}_{i\sigma_{y}^{(k)}}+\underbrace{i\sigma_{z}^{(k)}\sigma_{y}^{(k)}}_{\sigma_{x}^{(k)}})]\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}$
$=[\cos(\frac{\varphi}{2})\sigma_{+}^{(k)}-i\sin(\frac{\varphi}{2})\sigma_{+}^{(k)}]\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}=e^{-i\varphi/2}\sigma_{+}^{(k)}\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}$
$=e^{-i\varphi/2}\sigma_{+}^{(k)}U_{\varphi,k}^{\dagger}\underbrace{U_{\varphi,k}\prod_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{N}U_{\varphi,j}}_{U_{\varphi}}=e^{-i\varphi/2}(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})(\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}+i\sin(\frac{\varphi}{2})\sigma_{z}^{(k)})U_{\varphi}$
$=e^{-i\varphi/2}[\cos(\frac{\varphi}{2})(\sigma_{x}^{(k)}+i\sigma_{y}^{(k)})-i\sin(\frac{\varphi}{2})({i\sigma_{y}^{(k)}}+{\sigma_{x}^{(k)}})]U_{\varphi}=e^{-i\varphi}\sigma_{+}^{(k)}U_{\varphi}.$
One similarly finds $U_{\varphi}\sigma_{-}^{(k)}=e^{i\varphi}\sigma_{-}^{(k)}U_{\varphi}$. ∎
@@
$U_{\varphi,j}\sigma_{x}^{(j)}U_{\varphi,j}^{\dagger}$ $=(\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}-i\sin(\frac{\varphi}{2})\sigma_{z}^{(j)})\sigma_{x}^{(j)}(\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}+i\sin(\frac{\varphi}{2})\sigma_{z}^{(j)})$
$=(\cos(\frac{\varphi}{2})\mathbb{1}_{2^{N}}-i\sin(\frac{\varphi}{2})\sigma_{z}^{(j)})(\cos(\frac{\varphi}{2})\sigma_{x}^{(j)}+i\sin(\frac{\varphi}{2})\sigma_{x}^{(j)}\sigma_{z}^{(j)})$
$=\cos^{2}(\frac{\varphi}{2})\sigma_{x}^{(j)}+i\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{x}^{(j)}\sigma_{z}^{(j)}-i\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{z}^{(j)}\sigma_{x}^{(j)}+\sin^{2}(\frac{\varphi}{2})\sigma_{z}^{(j)}\sigma_{x}^{(j)}\sigma_{z}^{(j)})$
$=[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{x}^{(j)}+2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{y}^{(j)}.$
Similarly, one obtains $U_{\varphi,j}\sigma_{y}^{(j)}U_{\varphi,j}^{\dagger}=[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{y}^{(j)}-2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{x}^{(j)}$. Combining these results for $\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}$ and $\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}$, one obtains
$U_{\varphi}\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}U_{\varphi}^{\dagger}=$
=$=\left\{[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{x}^{(j)}+2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{y}^{(j)}\right\}\left\{[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{x}^{(j+1)}+2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{y}^{(j+1)}\right\},$
$U_{\varphi}\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}U_{\varphi}^{\dagger}=$
=$=\left\{[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{y}^{(j)}-2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{x}^{(j)}\right\}\left\{[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]\sigma_{y}^{(j+1)}-2\cos(\frac{\varphi}{2})\sin(\frac{\varphi}{2})\sigma_{x}^{(j+1)}\right\},$
which, using the identity $[\cos^{2}(\frac{\varphi}{2})-\sin^{2}(\frac{\varphi}{2})]^{2}=1-4\cos^{2}(\frac{\varphi}{2})\sin^{2}(\frac{\varphi}{2})$, leads to their sum $U_{\varphi}\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}U_{\varphi}^{\dagger}+U_{\varphi}\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}U_{\varphi}^{\dagger}=\sigma_{x}^{(j)}\sigma_{x}^{(j+1)}+\sigma_{y}^{(j)}\sigma_{y}^{(j+1)}$. ∎